# Effective Generative AI: The Human-Algorithm Centaur

### v1

@@ 4. Symbiotic Learning: A Core Characteristic of Centaurs
(5) $\hat{\theta}_{H}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{x\sim d_{human}}l_{human}(x;\theta).$
@@
(6) $\hat{\theta}_{H}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{x\sim d_{human}}l_{human}(x;\theta).$
@@
(7) $l_{human}(\hat{\theta}_{H})=\mathbb{E}_{(x,y1,y2)\sim d_{human}}[f_{1}(f_{2}(\hat{\theta}_{H}(x,y_{1})-\hat{\theta}_{H}(x,y_{2})))],$
@@
In some configurations, $\hat{\theta}_{H}$ is learnt first, and then used for learning $h_{\gamma}$. In other configurations, both $\hat{\theta}_{H}$ and $h_{\gamma}$ are jointly learned, where the output of $h_{{\theta}_{H}}$ is used as rewards to guide $h_{\gamma}$’s learning process and optimizing $l_{data}$ while at the same time outputs of $h_{\gamma}$ are used to further tune $\hat{\theta}_{H}$ using $l_{human}$. While human-guided rewards can be used for aligning different types of AI schemes, its most common use is for replacing or augmenting the reward function in reinforcement learning, creating Reinforcement Learning with Human Feedback (RLHF) schemes (rlhforiginal; intentions). As we will discuss later, RLHF in turn has played an important role in moving various generative AI models, including a vast majority of LLMs, from a traditional AI model to a centaur-based system by incorporating human intuition and symbiotic learning—the two distinguishing characteristics of centaurs.
@@
(9) $\hat{\theta}_{M}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{x\sim d_{data}}l_{data}(x;\theta),$
@@
(10) $\hat{\theta}_{symbiotic}=\operatorname*{arg\,min}_{\gamma\in\Gamma,\mathcal{C}(\gamma,\hat{\theta}_{M})\leq c_{1}}\mathbb{E}_{x\sim d_{human}}[l_{human}(x;\gamma;\hat{\theta}_{M})].$
@@
(11) $\hat{\theta}^{n}_{M}=\operatorname*{arg\,min}_{\theta\in\Theta_{n}}\mathbb{E}_{x\sim d_{data}}l_{data}(x;\theta),$
@@
(12) $\hat{\theta}^{k}_{H}=\operatorname*{arg\,min}_{\theta\in\Theta_{k}}\mathbb{E}_{x\sim d_{human}}l_{human}(x;\theta).$
@@
(13) $\hat{\theta}_{symbiotic}=\operatorname*{arg\,min}_{\gamma\in\Gamma}\mathbb{E}_{x\sim d_{human},d_{data}}[l_{comb}(x;\gamma)],$
@@ 5. Recent Advancements in LLMs and Their Cognitive Abilities: Evidence on the Power of Centaurs
Though LLMs have demonstrated extraordinary performance, the exact mechanism behind those impressive achievements is not entirely clear and has been the subject of an ongoing debate. Another, topic sparking ongoing debate is whether, and under which terms, LLMs can be used as cognitive models. A cognitive model is defined as “an approximation of one or more cognitive processes in humans or other animals for the purposes of comprehension and prediction” (cogdef). Thus, a model’s cognitive abilities can be captured either internally, by the extent to which the model’s internal learning and inference processes—captured via various neural activities—resemble those of humans; or externally, by the extent to which the model’s output follows human behaviors, decisions, and preferences. We claim that (a) modern LLMs already possess traits that highly resemble important human cognitive processes, and (b) a main reason for this is that they have moved from traditional AI models to centaurs by benefiting from direct incorporation of human intuition as well as symbiotic learning—the two main characteristics of centaurs. Building on these traits and enhancing them via more advanced symbiotic learning approaches or further incorporation of human intuition is an important building block for future endeavors in turning LLMs into much stronger cognitive models, capable of further representing human behavior.
@@
Enriching LLMs’ built-in, cognitive-like mechanisms such as attention and concept learning with symbiotic learning techniques and high-quality human-preference data has a great potential to further enrich LLMs’ cognitive abilities. This can be operationalized, for example, by using two-way in-context learning and human-preference-based attention weights as described above. Below we discuss two symbiotic-learning methods that have shown to be the underlying reasons why LLMs have gained cognitive abilities: human-preference-based supervised fine-tuning and human-guided rewards.
@@ 6. Discussion: Are Centaurs Always Superior?
In prior sections we have surveyed multiple techniques for creating centaurs. A key question that arises when considering the deployment of centaurs is the following: Does the fact that centaurs incorporate human intuition and symbiotic learning make them always have a better performance vis-a-vis some traditional AI models?
@@
However, ensuring that future centaurs perform well in both dimensions needs further research and deeper understandings. To illustrate this, we consider current efforts for increasing LLMs’ cognitive abilities as discussed in Section 5. Turning LLMs into cognitive models means that LLMs will also inherit at least some flaws of human cognition, not just its strengths. Indeed, it has been shown that on some tasks, LLMs’ ability to follow human’s intuitions, behaviors, or decisions negatively correlates with the model’s performance (studies2; bias1). That is to be expected as human decision-making process is oftentimes irrational, influenced by noise factors such as emotions and cognitive biases which might result in suboptimal decisions and behaviors (emotion). For instance, it has been shown in (bias1), (bias2), and (bias3) that LLMs share similar cognitive biases with humans, and that even though those biases lead to behaviors that are more similar to human behaviors, the biased behaviors also lead to deterioration in LLMs’ performance. While the latter has been recently highlighted in the context of LLMs, the cognitive bias problem presents a challenge in the design of any AI model where human biases are combined in the training process either explicitly by combining human preferences in the training algorithm, or implicitly by using biased training sets (fairness).
@@ 7. Conclusion: The Future of AI In Various Applications is Centaur
In this article, we presented the concept of centaurs as AI models that benefit from the paradigm of symbiotic learning as well as the incorporation of human intuition. We argued that centaurs have various advantages over traditional AI models. We also discussed that various human-level performances observed from recent LLMs including demonstration of cognitive ability stem from the fact that they have moved from traditional AI models to centaur-based systems.
@@
In closing, we note that while AI is still far from reaching “human-level intelligence”, the advancements in Generative AI have shown promising results for making use of centaurs in various applications. Centaurs might indeed be at the center of future AI developments. They might also be our main hope to get closer to reaching the goals that were stated by the founding fathers of AI. To realize this, it is useful to revisit what the founding fathers (John McCarthy, Marvin Minsky, Nathaniel Rochester, and Claude Shannon) stated in 1955 as part of their proposed 2-month, 10-man study of AI (to be held at Dartmouth):

### v4

@@ 5. Symbiotic Learning: A Core Characteristic of Centaurs
(5) $\hat{\theta}_{H}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{z\sim d_{human}}l_{human}(z;\theta).$
@@
(6) $\hat{\theta}_{H}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{z\sim d_{human}}l_{human}(z;\theta).$
@@
(7) $l_{human}(\hat{\theta}_{H})=\mathbb{E}_{(x,y_{1},y_{2})\sim d_{human}}[f_{1}(f_{2}(\hat{\theta}_{H}(x,y_{1})-\hat{\theta}_{H}(x,y_{2})))],$
@@
In some configurations, $\hat{\theta}_{H}$ is learnt first, and then used for learning $h_{\gamma}$. In other configurations, both $\hat{\theta}_{H}$ and $h_{\gamma}$ are jointly learned, where the output of $h_{{\theta}_{H}}$ is used as rewards to guide $h_{\gamma}$’s learning process and optimizing $l_{data}$ while at the same time outputs of $h_{\gamma}$ are used to further tune $\hat{\theta}_{H}$ using $l_{human}$. While human-guided rewards can be used for aligning different types of AI schemes, its most common use is for replacing or augmenting the reward function in reinforcement learning, creating Reinforcement Learning with Human Feedback (RLHF) schemes (rlhforiginal; intentions). As we will discuss later, RLHF in turn has played an important role in moving various generative AI models, including a vast majority of LLMs, from a pure AI model to a centaur-based system by incorporating human intuition and symbiotic learning—the two distinguishing characteristics of centaurs.
@@
(9) $\hat{\theta}_{M}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{z\sim d_{data}}l_{data}(z;\theta),$
@@
(10) $\hat{\theta}_{symbiotic}=\operatorname*{arg\,min}_{\gamma\in\Gamma,\mathcal{C}(\gamma,\hat{\theta}_{M})\leq c_{1}}\mathbb{E}_{z\sim d_{human}}[l_{human}(z;\gamma;\hat{\theta}_{M})].$
@@
(11) $\hat{\theta}^{n}_{M}=\operatorname*{arg\,min}_{\theta\in\Theta_{n}}\mathbb{E}_{z\sim d_{data}}l_{data}(z;\theta),$
@@
(12) $\hat{\theta}^{k}_{H}=\operatorname*{arg\,min}_{\theta\in\Theta_{k}}\mathbb{E}_{z\sim d_{human}}l_{human}(z;\theta).$
@@
(13) $\hat{\theta}_{symbiotic}=\operatorname*{arg\,min}_{\gamma\in\Gamma}\mathbb{E}_{z\sim d_{human},d_{data}}[l_{comb}(z;\gamma)],$
@@ 6. Recent Advancements in LLMs and Their Cognitive Abilities: Evidence on the Power of Centaurs
Though LLMs have demonstrated extraordinary performance, the exact mechanism behind those impressive achievements is not entirely clear and has been the subject of an ongoing debate. Another, topic sparking ongoing debate is whether, and under which terms, LLMs can be used as cognitive models. A cognitive model is defined as “an approximation of one or more cognitive processes in humans or other animals for the purposes of comprehension and prediction” (cogdef). Thus, a model’s cognitive abilities can be captured either internally, by the extent to which the model’s internal learning and inference processes—captured via various neural activities—resemble those of humans; or externally, by the extent to which the model’s output follows human behaviors, decisions, and preferences. We claim that (a) modern LLMs already possess traits that highly resemble important human cognitive processes, and (b) a main reason for this is that they have moved from pure AI models to centaurs by benefiting from direct incorporation of human intuition as well as symbiotic learning—the two main characteristics of centaurs. Building on these traits and enhancing them via more advanced symbiotic learning approaches or further incorporation of human intuition is an important building block for future endeavors in turning LLMs into much stronger cognitive models, capable of further representing human behavior.
@@
When faced with complex problems, one reasoning strategy employed by humans’ cognition is “divide-and-concur:" decomposing the problem into manageable, intermediate steps and solving each one before giving the final answer (wei2022chain). Such intermediate sub-problems can be seen as a series of logical deductions that sequentially lead to a conclusive answer. The benefits of such sequential decomposition reasoning in humans raised the question of whether such “intermediate thinking" patterns can improve LLMs’ reasoning abilities, leading to the development of Chain-of-Thought reasoning (wei2022chain). In its simplest form, Chain-of-Thought (CoT) denotes a series of intermediate natural language reasoning steps that lead to a final output (reasoning).
Until recently, CoT has been incorporated within LLMs primarily using CoT Prompting: instead of demonstrations solely composed of a (query, answer) pair, CoT Prompting additionally includes a rationale for each example, encouraging the model to verbalize the intermediate reasoning steps for solving the task (reasoning). Newer models, such as the latest release of OpenAI’s o1 (o12), take one step forward and incorporate CoTs within the model’s learning and reasoning process. In this way, human input in the form of CoTs is explicitly combined within the model’s architecture rather than being queried in an ad-hoc fashion by the end user as part of the prompt. The o1 model is trained to follow human-produced CoTs by using an RL scheme designed to improve the model’s alignment not only on final answers but also on their corresponding CoTs—the reasoning process that led to the final answers (planning; o11).
While CoT reasoning can certainly be seen as one step forward towards creating centaur models, as mentioned earlier, such models are still different than centaurs. A critical difference is in that they include only conscious reasoning patterns, depriving LLMs from any training data representing unconscious reasoning patterns. Thus, models trained on CoT human input alone may not be considered as centaurs: they lack true human intuition, at least based on the definition of intuition we discussed earlier.
Indeed, recent works show that while o1 significantly outperforms GPT-4 on problem-solving tasks such as math and coding problems, it performs on par (and sometimes, subpar) with GPT-4 on ambiguous tasks, highly-abstract tasks, tasks requiring flexibility an adaptability to real-time dynamics, or ones that require knowledge gained through experience (o11; o12). Furthermore, the iterative reasoning process of o1 makes it significantly slower than GPT-4, leading to longer output times (o11).
Nontheless, enriching LLMs’ built-in, cognitive-like mechanisms such as attention and concept learning with symbiotic learning techniques and high-quality human-preference data has a great potential to further enrich LLMs’ cognitive abilities. This can be operationalized, for example, by using two-way in-context learning and human-preference-based attention weights as described above. Below we discuss two specific symbiotic-learning methods that have shown to be the underlying reasons why LLMs have gained cognitive abilities: human-preference-based supervised fine-tuning and human-guided rewards.
@@ 7. Discussion: Are Centaurs Always Superior?
In prior sections we have surveyed multiple techniques for creating centaurs. A key question that arises when considering the deployment of centaurs is the following: Does the fact that centaurs incorporate human intuition and symbiotic learning make them always have a better performance vis-a-vis some pure AI models?
@@
However, ensuring that future centaurs perform well in both dimensions needs further research and deeper understandings. To illustrate this, we consider current efforts for increasing LLMs’ cognitive abilities as discussed in Section 6. Turning LLMs into cognitive models means that LLMs will also inherit at least some flaws of human cognition, not just its strengths. Indeed, it has been shown that on some tasks, LLMs’ ability to follow human’s intuitions, behaviors, or decisions negatively correlates with the model’s performance (studies2; bias1). That is to be expected as human decision-making process is oftentimes irrational, influenced by noise factors such as emotions and cognitive biases which might result in suboptimal decisions and behaviors (emotion). For instance, it has been shown in (bias1), (bias2), and (bias3) that LLMs share similar cognitive biases with humans, and that even though those biases lead to behaviors that are more similar to human behaviors, the biased behaviors also lead to deterioration in LLMs’ performance. While the latter has been recently highlighted in the context of LLMs, the cognitive bias problem presents a challenge in the design of any AI model where human biases are combined in the training process either explicitly by combining human preferences in the training algorithm, or implicitly by using biased training sets (fairness). Another potential threat in developing centaur-type models driven by human intuition is the development of dark-triad-like traits (darktriad) within LLMs. At scale, such traits may pose serious safety risks by driving the model to engage is an array of negative behaviors from systematic deception and content manipulation to power-seeking behaviors, potentially with the collaboration of other centaur-like agents (risk1; risk2; risk3).
@@ 8. Conclusion: The Future of AI In Various Applications is Centaur
In this article, we presented the concept of centaurs as AI models that benefit from the paradigm of symbiotic learning as well as the incorporation of human intuition. We argued that centaurs have various advantages over other AI models. We also discussed that various human-level performances observed from recent LLMs including demonstration of cognitive ability stem from the fact that they have moved from pure AI models to centaur-based systems.
@@
In closing, we note that while AI is still far from reaching “human-level intelligence,” the advancements in Generative AI have shown promising results for making use of centaurs in various applications. Centaurs might indeed be at the center of future AI developments. They might also be our main hope to get closer to reaching the goals that were stated by the founding fathers of AI. To realize this, it is useful to revisit what the founding fathers (John McCarthy, Marvin Minsky, Nathaniel Rochester, and Claude Shannon) stated in 1955 as part of their proposed 2-month, 10-man study of AI (to be held at Dartmouth):